The return on US T-Bills is 6%, inflation is 3% and the risk premium of the S&P 500 is 129, If a stock has a beta of 1.3 , what is the expected rearrateofreturn forthe stock? Hhon A. 13.8% B. 21.6% C. 18.1% D. 10.5% 1.0%

the answer is C. Please show the process.
Answer #1

Expected return=Risk free rate+Risk premium*Beta

=6+(12*1.3)=21.6%

Expected real rate=(1+nominal rate)/(1+inflation rate)-1

=(1+0.216)/(1+0.03)-1

=(1.216/1.03)-1

=18.1%(Approx).